Leyton Lumber Company has sales of $9 million per year, all on credit terms calling for payment within 30 days, and its accounts receivable are $1.35 million. Assume 365 days in year for your calculations.

  1. What is Leyton's DSO? Round your answer to two decimal places.
      days

  2. What would DSO be if all customers paid on time? Round your answer to two decimal places.
      days

  3. How much capital would be released if Leyton could take actions that led to on-time payments? Round your answer to the nearest cent. Write out your answer completely. For Example, 13.2 million should be entered as 13,200,000.
Answer #1

DSO = Receivables / Sales * 365

= 1.35 / 9 * 365

= 55 days

If all the customers paid on time, the DSO would be the credit period offered by Leyton Lumber Co. which is = 30 days in this case

Calculate the accounts receivable when the DSO = 30 days

Receivables = DSO * Sales / 365

= 30 * 9000000 /365

= 739726.03

Difference in receivables = 1350000 - 739726.03

= 610273.97

Hence if the company takes action to bring timely payments, the receivables would reduce by $610274 and this amount would be released.